# A Survey of Reinforcement Learning For Economics

### 15 Mar 2026 v2

@@ 13.6 Simulation Study: Confounded Retail Pricing MDP
The action affects the next state only through the mediator: $P(s{+}1\mid s,M_{t},Z_{t})$ depends on $M_{t}$ and $Z_{t}$ but not on $A_{t}$ or $U_{t}$ directly. This enables all four identification strategies simultaneously: $Z_{t}$ satisfies the backdoor criterion, $M_{t}$ satisfies the front-door criterion, $\text{IV}_{t}$ satisfies relevance and exclusion, and $W_{t}^{(1)},W_{t}^{(2)}$ satisfy the proximal conditions.207207207Rewards are r(s,a)=1r(s,a)=-1 for $s<4$, $\gamma=0.9$. The target policy always promotes. The true interventional transition probability is $P(s{+}1\mid s,\operatorname{do}(\text{promote}))=0.615$.
@@
I compare six estimators: oracle, naive (biased per Lemma L1), backdoor (Equation 111), front-door (Equation 112), Wald IV (Equation 114), and proximal (Equations 115116).208208208Each configuration uses 2,000 trajectories averaged over 20 seeds.
@@
Table 26 reports bias and RMSE across confounding strengths. The naive estimator’s bias grows monotonically with $\rho$ because observational transitions overestimate promotion success: the promote action is more likely when $U_{t}=1$, which correlates with favorable conditions, so $\hat{P}_{\text{obs}}$ exceeds the true interventional probability, and this per-step bias compounds through the Bellman recursion. The backdoor and front-door estimators eliminate bias at all $\rho$, validating Theorem 8 and Equation (112). The IV estimator maintains low bias but higher variance due to the Wald ratio’s sensitivity to instrument strength; panel (b) of Figure 27 illustrates this classic bias-variance tradeoff. The proximal estimator achieves low bias with moderate variance, confirming that bridge functions recover causal effects from noisy proxies.
Table 26: Unified identification in a confounded retail pricing MDP: bias and RMSE of five estimators across confounding strengths $\rho$ (20 seeds, 2,000 trajectories per seed).
Naive Backdoor Front-door IV Proximal
$\rho$ Bias (SE) RMSE Bias (SE) RMSE Bias (SE) RMSE Bias (SE) RMSE Bias (SE) RMSE
0.0 0.010-0.010 ($0.005$) $0.024$ 0.008-0.008 ($0.005$) $0.024$ 0.001-0.001 ($0.005$) $0.023$ 0.006-0.006 ($0.028$) $0.122$ 0.011-0.011 ($0.006$) $0.028$
0.2 +0.060+0.060 ($0.005$) $0.064$ 0.007-0.007 ($0.006$) $0.025$ +0.003+0.003 ($0.006$) $0.025$ +0.012+0.012 ($0.028$) $0.125$ 0.012-0.012 ($0.006$) $0.027$
0.4 +0.127+0.127 ($0.005$) $0.129$ 0.008-0.008 ($0.006$) $0.026$ +0.001+0.001 ($0.006$) $0.027$ +0.011+0.011 ($0.029$) $0.127$ 0.015-0.015 ($0.006$) $0.028$
0.6 +0.194+0.194 ($0.005$) $0.195$ 0.007-0.007 ($0.006$) $0.028$ +0.001+0.001 ($0.006$) $0.028$ +0.020+0.020 ($0.030$) $0.131$ 0.014-0.014 ($0.007$) $0.033$
0.8 +0.258+0.258 ($0.005$) $0.259$ 0.007-0.007 ($0.006$) $0.029$ 0.004-0.004 ($0.007$) $0.030$ +0.032+0.032 ($0.030$) $0.136$ 0.015-0.015 ($0.007$) $0.034$
1.0 +0.321+0.321 ($0.005$) $0.322$ 0.005-0.005 ($0.006$) $0.028$ 0.004-0.004 ($0.007$) $0.030$ +0.055+0.055 ($0.025$) $0.123$ 0.015-0.015 ($0.008$) $0.037$
@@
Figure 27: (a) Bias of five OPE estimators relative to the oracle as a function of confounding strength $\rho$. (b) IV estimator bias distribution vs. instrument strength at $\rho=1$. Dashed red line: naive estimator bias.
14 Conclusion
Economics has long relied on dynamic programming to formalize and solve problems of sequential decision-making. Yet, for decades, the complexity of our models has been dictated less by the economic phenomena we wish to study and more by the computational limits of the curse of dimensionality.209209209Rust (2019) asks the same question of dynamic programming itself, surveying its practical impact across domains and identifying the gap between theoretical power and real-world deployment. Reinforcement learning offers one pragmatic path forward. By replacing exhaustive state-space enumeration with sampling and function approximation, RL trades the absolute convergence guarantees of classical dynamic programming for the capacity to operate in high-dimensional, continuous, and complex strategic environments.
As this survey has sought to illustrate, RL is not a departure from dynamic programming, but a scalable and natural extension of it. It does come with its own issues. Deep reinforcement learning, in particular, remains brittle, sensitive to hyperparameters, and vulnerable to the structural instabilities of the deadly triad. Furthermore, its success in structural estimation and policy design relies heavily on computational tricks and simplifying assumptions, the availability of highly accurate simulators, and clearly defined transition or reward functions.
However, the intersection of these two fields is deeply complementary. While RL equips economics with new tools, economics provides RL with much-needed structure. As demonstrated in applications ranging from dynamic pricing to causal inference, incorporating fundamental economic concepts, such as revealed preference, parametric assumptions, or causal identification, can tame RL’s sample complexity and prevent algorithms from converging on systematically biased policies.
Looking ahead, we should view reinforcement learning simply as another tool in the applied economist’s toolkit. As these algorithms mature and their theoretical foundations solidify, the novelty of using ”AI” to solve economic models will likely fade. However, one can expect interesting cross fertilization between these computational fields, that gradually expands the frontier of the models we can solve, the mechanisms we can design, and the behaviors we can understand.

### 17 Mar 2026 v3

@@ 12.6 Simulation Study: Confounded Retail Pricing MDP
The action affects the next state only through the mediator: $P(s{+}1\mid s,M_{t},Z_{t})$ depends on $M_{t}$ and $Z_{t}$ but not on $A_{t}$ or $U_{t}$ directly. This enables all four identification strategies simultaneously: $Z_{t}$ satisfies the backdoor criterion, $M_{t}$ satisfies the front-door criterion, $\text{IV}_{t}$ satisfies relevance and exclusion, and $W_{t}^{(1)},W_{t}^{(2)}$ satisfy the proximal conditions.224224224Rewards are r(s,a)=1r(s,a)=-1 for $s<4$, $\gamma=0.9$. The target policy always promotes. The true interventional transition probability is $P(s{+}1\mid s,\operatorname{do}(\text{promote}))=0.615$.
@@
I compare six estimators: oracle, naive (biased per Lemma L1), backdoor (Equation 119), front-door (Equation 120), Wald IV (Equation 122), and proximal (Equations 123124).225225225Each configuration uses 2,000 trajectories averaged over 20 seeds.
@@
Figure 27 reports bias and RMSE across confounding strengths (20 seeds, 2,000 trajectories per seed). The naive estimator’s bias grows monotonically with $\rho$ because observational transitions overestimate promotion success: the promote action is more likely when $U_{t}=1$, which correlates with favorable conditions, so $\hat{P}_{\text{obs}}$ exceeds the true interventional probability, and this per-step bias compounds through the Bellman recursion. The backdoor and front-door estimators eliminate bias at all $\rho$, validating Theorem 8 and Equation (120). The IV estimator maintains low bias but higher variance due to the Wald ratio’s sensitivity to instrument strength; panel (c) illustrates this classic bias-variance tradeoff. The proximal estimator achieves low bias with moderate variance, confirming that bridge functions recover causal effects from noisy proxies.
@@
Figure 27: (a) Bias of five OPE estimators as a function of confounding strength $\rho$. (b) RMSE of five estimators as a function of $\rho$. (c) IV estimator bias distribution vs. instrument strength at $\rho=1$; dashed red line is naive estimator bias.
13 Discussion
This concluding section develops concrete research agendas at the intersection of reinforcement learning and economics as well as lists the bottlenecks and open challenges.
13.1 How Economics Improves Reinforcement Learning
Reinforcement learning’s most celebrated successes, from Atari to Go to protein folding, share a common ingredient: a cheap, fast, and accurate simulator that generates unlimited training data. Economics largely lacks this ingredient. Every application in Section 7 demanded a custom environment, and the engineering cost of building these environments often dominated the cost of training the RL agent itself. Economics is, however, uniquely positioned to fill this gap. Economic models encode variable selection, causal structure, and institutional constraints that determine how agents respond to interventions. The Lucas critique warns that correlational simulators, trained on observational data without structural assumptions, will break under policy changes (Section 12). Building economic simulators that respect causal identification and correctly model agent responses to rule changes is an important problem.
A recurring theme across the survey is that economic structure, when available, can dramatically reduce the sample complexity of learning. The knowledge ladder in Section 10 shows that imposing demand structure on a pricing problem can reduce cumulative regret from $\Theta(T)$ to $O(\log T)$. The pattern extends beyond bandits: structural assumptions yield similar gains in dynamic estimation (Section 8) and preference learning (Section 11). These reflect the difference between learning in an unstructured space and learning in one shaped by economic theory. Formalizing this intuition, identifying which structural assumptions yield which complexity reductions and under what conditions, is an active research frontier.
The RLHF and DPO frameworks studied in Section 11 rely on the Bradley-Terry model, one of the simplest members of the discrete choice family. Economics offers a rich toolkit for moving beyond it. Mixed logit models accommodate heterogeneous preferences across annotators. Revealed preference theory provides axiomatic consistency constraints, such as GARP and stochastic transitivity, that can serve as regularizers on learned reward models. The preference learning simulation in Section 11 illustrates the cost of misspecification (Figure 23). Integrating econometric tools for preference elicitation, model selection, and specification testing into the RLHF pipeline is a natural direction.
13.2 How Reinforcement Learning Advances Economics
Dynamic programming has always offered economics a prescriptive capability: given a model, compute the optimal policy. In practice, this capability has been limited by the curse of dimensionality to models with small, discrete state spaces. RL relaxes this constraint. TD-based methods make structural models tractable at state-space scales where NFXP is infeasible (Section 8), and real-world deployments confirm that RL can compute policies of practical value (Section 7). These results suggest a prescriptive role for RL in economics: not merely estimating model parameters (the traditional econometric task), but computing the policies those parameters imply.
Social science and policy research work with vast quantities of observational data from settings where controlled experimentation is impossible or unethical. Offline RL, which learns policies from logged data without further interaction, is a natural fit. The simulation in Section 11 illustrates both the promise and the limits: algorithms with distributional shift correction exceed the behavioral baseline, while those without it degrade below it, and performance depends critically on data support (Table 20, Figure 20). This means that offline RL is not a black box that extracts optimal policies from any observational dataset. It is a framework with clearly delineated conditions under which learning is possible, conditions that connect directly to familiar econometric concepts like overlap and common support. Developing standardized benchmarks and digital twins for economic offline RL can enable firms and policy-makers to design optimal policies from data generated under old ones.
RL also offers a descriptive model of how boundedly rational agents learn in economic environments. The simulations in Section 9 demonstrate that independent Q-learning agents converge to Nash equilibrium through trial and error, providing “as-if” microfoundations for equilibrium concepts: the equilibrium arises not from common knowledge of rationality, but from a simple adaptive process.
When RL-trained agents are actually deployed in markets, they become economic actors subject to empirical scrutiny. Algorithmic pricing agents, for instance, have been shown to sustain supra-competitive prices through reward-based learning without explicit communication. This is not a speculative concern; competition authorities in the EU, US, and UK are actively investigating whether algorithmic coordination constitutes tacit collusion.226226226The companion thesis (Rawat, 2026) develops a framework for evaluating algorithmic inefficiency and collusion risk in algorithmically mediated markets, combining simulators with factorial experimental designs. As algorithmic agents proliferate in pricing, trading, content recommendation, and resource allocation, the empirical study of machine behavior using economic tools becomes an increasingly important research program.
13.3 Open Challenges
For researchers accustomed to estimators with well-characterized asymptotic properties, the deadly triad represents a significant barrier to adoption. Deep RL algorithms exhibit seed sensitivity, overestimation cascades, and plasticity loss that make reproducibility difficult even in controlled settings (Section 6).
Multi-agent RL faces fundamental problems (Section 9). Computing Nash equilibria is PPAD-complete, and RL does not escape this hardness. In economic games with multiple equilibria, agents performing Bellman backups may select different equilibria for different states, causing value iteration to cycle or diverge. The literature on equilibrium refinement, focal points, and coordination mechanisms may offer partial resolutions, but a general solution remains open.
The absence of standardized economic simulators is part of a broader infrastructure gap. Industrial RL deployments require dedicated engineering teams, GPU clusters, and months of hyperparameter tuning (Section 7, Section 6). Reducing these barriers through shared simulation environments and accessible software libraries would accelerate adoption.
13.4 Conclusion
Reinforcement learning is a welcome addition to the economist’s toolkit but is not a replacement for existing methods. It extends the reach of dynamic programming to problems that were previously intractable, and it connects naturally to econometric frameworks through shared mathematical foundations. The research agendas outlined above, from building economic simulators to formalizing the role of structural assumptions in sample complexity to developing inference procedures for learned policies, are concrete and tractable. Progress on them will require sustained collaboration between the two fields, drawing on economics for “structure” and institutional context and on RL for computation.
Appendix A Glossary of Acronyms and Terms
Acronyms
A2C/A3C Advantage Actor-Critic / Asynchronous Advantage Actor-Critic. Policy gradient algorithms that use an advantage function baseline to reduce variance (Section 4).
BCQ Batch-Constrained Q-learning. An offline RL algorithm that restricts the learned policy to actions supported by the behavioral data (Section 11.2).
BwK Bandits with Knapsacks. A bandit framework with resource constraints, where an agent must balance exploration and exploitation subject to a budget (Section 10).
CCP Conditional Choice Probabilities. Choice probabilities conditional on state, used in structural estimation as an alternative to full-solution methods (Section 8).
CFR Counterfactual Regret Minimization. An iterative algorithm for computing Nash equilibria in extensive-form games by minimizing regret at each information set (Section 9).
CQL Conservative Q-Learning. An offline RL algorithm that adds a penalty for overestimating Q-values on out-of-distribution actions (Section 11.2).
DDC Dynamic Discrete Choice. A class of structural econometric models in which agents make sequential discrete decisions under uncertainty (Section 8).
DDPG Deep Deterministic Policy Gradient. An actor-critic algorithm for continuous action spaces that combines a deterministic policy gradient with a learned Q-function (Section 4).
DP Dynamic Programming. A collection of algorithms that compute optimal policies by solving the Bellman equation, given a known model of the environment (Section 4).
DPO Direct Preference Optimization. A method that optimizes a language model directly on preference data without training a separate reward model (Section 11.4).
DQN Deep Q-Network. Q-learning with a neural network function approximator, target network, and experience replay (Section 4).
ETC Explore-Then-Commit. A bandit algorithm that explores uniformly for a fixed number of rounds, then commits to the empirically best arm (Section 10).
FQI/FVI Fitted Q-Iteration / Fitted Value Iteration. Approximate dynamic programming algorithms that use supervised learning to fit value functions from batch data (Section 5.2.5).
GLIE Greedy in the Limit with Infinite Exploration. A condition on exploration schedules ensuring convergence to the optimal policy (Section 4).
GMM Generalized Method of Moments. An econometric estimation method that matches sample moments to population moment conditions (Section 8).
IPW Inverse Probability Weighting. A method for correcting distributional mismatch by reweighting observations by the inverse of their selection probability (Section 12).
IQL Implicit Q-Learning. An offline RL algorithm that avoids querying out-of-distribution actions by using expectile regression on the value function (Section 11.2).
IRL Inverse Reinforcement Learning. The problem of inferring a reward function from observed behavior, assuming the agent acts approximately optimally (Section 4).
IV Instrumental Variables. An econometric technique for identifying causal effects in the presence of endogeneity by exploiting exogenous variation (Section 12).
KL Kullback-Leibler divergence. A measure of how one probability distribution diverges from a reference distribution, used in trust-region and regularization methods (Section 4).
LLM Large Language Model. A neural network trained on large text corpora, the primary object of alignment via RLHF and DPO (Section 11.4).
MARL Multi-Agent Reinforcement Learning. RL in settings with multiple interacting agents, where each agent’s optimal policy depends on the others’ behavior (Section 9).
MC Monte Carlo. Methods that estimate value functions from complete episode returns rather than bootstrapped estimates (Section 4).
MDP Markov Decision Process. A formal model of sequential decision-making defined by states, actions, transition probabilities, rewards, and a discount factor (Section 4).
MLE Maximum Likelihood Estimation. A statistical method that estimates parameters by maximizing the likelihood of the observed data (Section 8).
NE Nash Equilibrium. A strategy profile in which no player can improve their payoff by unilaterally deviating (Section 9).
NFXP Nested Fixed Point. Rust’s algorithm for structural estimation of DDC models that nests value function iteration inside a maximum likelihood loop (Section 8).
NPG Natural Policy Gradient. A policy gradient method that preconditions the gradient by the inverse Fisher information matrix (Section 4).
OPE Off-Policy Evaluation. Estimating the expected return of a target policy using data generated by a different behavioral policy (Section 12).
PEVI Pessimistic Value Iteration. An offline RL algorithm that subtracts an uncertainty penalty from value estimates to avoid overestimation on unseen state-action pairs (Section 12).
PI Policy Iteration. A dynamic programming algorithm that alternates between policy evaluation and policy improvement until convergence (Section 4).
POMDP Partially Observable MDP. An MDP in which the agent cannot directly observe the state and must maintain beliefs from noisy observations (Section 12).
PPO Proximal Policy Optimization. A policy gradient algorithm that constrains updates to a trust region via a clipped surrogate objective (Section 4).
RLHF Reinforcement Learning from Human Feedback. A framework for aligning model behavior with human preferences by training a reward model from pairwise comparisons and optimizing a policy against it (Section 11.4).
SAC Soft Actor-Critic. An off-policy actor-critic algorithm that maximizes a combined objective of expected return and policy entropy (Section 4).
SARSA State-Action-Reward-State-Action. An on-policy TD control algorithm that updates Q-values using the action actually taken in the next state (Section 4).
SFT Supervised Fine-Tuning. The initial stage of LLM alignment in which the model is trained on curated demonstrations before preference optimization (Section 11.4).
SPE Subgame Perfect Equilibrium. A refinement of Nash equilibrium requiring that strategies constitute a Nash equilibrium in every subgame (Section 9).
TD Temporal Difference. A class of prediction and control algorithms that update value estimates using bootstrapped targets rather than complete returns (Section 5.2).
TRPO Trust Region Policy Optimization. A policy gradient algorithm that constrains each update to lie within a KL-divergence trust region of the previous policy (Section 4).
TS Thompson Sampling. A Bayesian bandit algorithm that selects actions by sampling from the posterior distribution over reward parameters (Section 10).
UCB Upper Confidence Bound. A bandit algorithm that selects the arm with the highest sum of empirical mean and an exploration bonus (Section 10).
VI Value Iteration. A dynamic programming algorithm that iteratively applies the Bellman optimality operator until the value function converges (Section 4).